# November, 2025

### v1

@@ 3.2 Attractive force by potential
In the next section, we will discuss a genuine “black hole hydrogen atom,” namely, a negatively charged particle trapped by a positively charged black hole. Before discussing this truly ideal situation, we first consider a simpler setup, i.e., a particle subject to a hypothetical attractive potential in the Schwarzschild geometry. We assume that there is the potential $V=-\frac{e^{2}}{r}$ producing an electric-like attractive force. We here employ the unit $4\pi\varepsilon_{0}=1$ and $e$ is the charge of the particle. We call this electron in the following. This setup gives a simple model for the hydrogen atom near the black hole. The equation in the radial direction is given by
@@
We find that $r=0$ and $r=a$ are the regular singular points. At the horizon $r=a$, the solution behaves like $R(r)\sim(r-a)^{0}$ and $(r-a)^{\frac{1}{2}}$. Then $R(r)$ is regular but $R^{\prime}(r)$ generically diverges at $r=a$.
Since the geometry is asymptotically flat, the equation coincides with the ordinary one for the hydrogen atom in $r\to\infty$. In this regime the solution should coincides with the well-known wavefunction. Since the energy spectrum of the hydrogen atom is determined by the asymptotic boundary condition, we can employ the energy spectrum $E_{n}=-\frac{\mu e^{4}}{2n^{2}\hbar^{2}}$ ($n=1,2,3,\ldots$) for the flat space even in the presence of the black hole. Following the prescription presented in the previous subsection, we find that the equation for the radial direction $R(x)=\mathrm{e}^{i\sigma x}H(x)$ is given by
$H^{\prime\prime}(x)+\left\{2\sigma+\frac{\frac{3}{2}}{x}-\frac{\frac{1}{2}}{x-1}\right\}H^{\prime}(x)+\frac{1}{x(x-1)}\left\{(2\sigma-a^{2}\kappa^{2}+as^{2})x-\frac{3}{2}\sigma-l(l+1)\right\}H(x)=0,$
where we have again defined $x=\frac{r}{a}$, $\sigma=\pm a\kappa$, $\kappa^{2}=\frac{-2\mu E_{n}}{\hbar^{2}}$ and $s^{2}=\frac{2\mu e^{2}}{\hbar^{2}}$. The general solution is therefore given by the confluent Heun’s function $H(\gamma,\delta,\varepsilon,\alpha,q,x)$ where the parameters are given by $\gamma=\frac{3}{2}$, $\delta=-\frac{1}{2}$, $\varepsilon=2\sigma$, $\alpha=2\sigma-a^{2}\kappa^{2}+as^{2}$ and $q=-\frac{3}{2}\sigma-l(l+1)$.
We find that it is convenient to visualize the wavefunction by solving the equation numerically. To this end, we rewrite the equation (3.15) into the following form:
@@
Figure 3: The radial probability densities of the wavefunction $\sqrt{g}|R(r)|^{2}r^{2}$. Comparison between the hydrogen atom in the flat space and in the black hole background cases for $a=1$. The figure corresponds to the $3p$ orbital ($n=3$, $l=1$) of the hydrogen atom. We will always use the parameters $\mu=e=\hbar=1$ in the following.
Figure 4: The 3D plot of the probability density $|\psi|^{2}$ for the particle under the attractive force. The flat space (left) and the black hole background (right) cases. The figures correspond to the $3d$ orbital ($n=3$, $l=2$) of the hydrogen atom. The black hole horizon $a=5$ is shown in the black color.
Even though the wavefunction is strongly concentrated near the horizon, the peak appears to be displaced to outside when it slightly farther away from the horizon. This behavior can be interpreted as follows. Although electrons are gravitationally attracted toward the black hole and extremely localized in a thin shell near the horizon, the uncertainty principle causes their kinetic energy to increase when they are too close to the horizon surface. Interestingly, the localization of electrons due to the strong gravitational attraction of the black hole has the opposite effect of repulsing particles away from it. These two forces balance to minimize the energy, resulting in a finite spread and peak positions of the wavefunction. Paradoxically, the stronger the gravitational attraction toward the center, the farther the electrons are pushed away. The difference between the bound state and the free particle discussed in Sec. 3.1 is that the former makes standing waves and localize sharply, while the latter does not. Then it is natural that the former is more subject to the uncertainty relation.
However when the black hole becomes sufficiently large, the situation changes. Fig. 5 shows a comparison of the radial probability densities obtained by varying the horizon position $r=a$.
(a) $a=1$.
(b) $a=4$.
(c) $a=8$.
(d) $a=14$.
Figure 5: Comparison of the radial probability density $\sqrt{g}|R(r)|^{2}r^{2}$ for different $a$ values in the Schwarzschild background with the potential $V$. Each figure corresponds to the $3p$ orbital ($n=3$, $l=1$) of the hydrogen atom. The red dotted line denotes the event horizon.
When the horizon position changes, the peak positions of the wavefunction moves accordingly. By changing the horizon position and making it larger than the peak position of the probability density, the corresponding peak appears to be absorbed into the horizon. Then gravity eventually prevails over the quantum effects and electrons will inevitably fall into the black hole.
@@ 4 Black hole atoms
We now study the genuine “black hole atoms.” We introduce the single non-extremal Reissner–Nordström black hole as the positively charged atomic nucleus. In the following, we examine the quantum wavefunction for a negatively charge particle trapped by the black hole. This is nothing but the hydrogen atom composed of a black hole nucleus and an electron.
@@ 4.1 Black hole hydrogen atom
The non-extremal Reissner–Nordström black hole metric in the Cartesian coordinate is given by
@@
where $r^{2}=x^{2}+y^{2}+z^{2}$ and $a=\frac{2G_{\mathrm{N}}M}{c^{2}}$, $Q^{2}=\frac{G_{\mathrm{N}}e^{2}}{c^{4}}$ are constants. The outer and the inner horizons are at $r=r_{+}$ and $r=r_{-}$, respectively, where we have defined $r_{\pm}=\frac{a\pm\sqrt{a^{2}-4Q^{2}}}{2}$. The gauge field $A_{0}=-\frac{1}{c}\frac{e}{r}$ in the solution provides the electrostatic potential. The time slice of the solution (4.1) gives a three-dimensional spatial curved space. The spacial metric is given by
@@
We have also $g=\det g_{ij}=\frac{r^{2}}{(r-r_{+})(r-r_{-})}$. Then, after some calculations, the Schrödinger equation in the $(r,\theta,\phi)$ system is given by
@@
By substituting the decomposition $\psi(r,\theta,\phi)=R(r)Y(\theta,\phi)$ in the equation, we find that the angular function $Y(\theta,\phi)$ satisfies the same equation in the Schwarzschild case and hence it is the spherical harmonics. On the other hand, the equation in the radial direction is given by
@@
Therefore the $R(r)$ is regular at the horizons but its derivative diverges as in the case of the Schwarzschild black hole. Now we look for bound states where $E<0$. The asymptotic consistency again results in $E_{n}=-\frac{\mu e^{4}}{2n^{2}\hbar^{2}}$. Using the same parameters as discussed before, the equation (4.13) is rewritten as
$\begin{aligned} \displaystyle&(\rho^{2}-2\kappa a\,\rho+4\kappa^{2}Q^{2})\,\frac{\partial^{2}u}{\partial\rho^{2}}+\Bigl\{-\rho^{2}+2\rho\,(l+1+\kappa a)-\kappa a\,(4l+3)+\frac{4\kappa^{2}Q^{2}}{\rho}\,(2l+1-\rho)\Bigr\}\frac{\partial u}{\partial\rho}\\ &+\Bigg[\rho\Bigl(n-l-1-\frac{\kappa a}{2}\Bigr)+\kappa a\Bigl(2l+\frac{3}{2}\Bigr)-\left\{\frac{\kappa a\,l}{\rho}+\frac{2\kappa^{2}Q^{2}}{\rho}\right\}(2l+1)+\frac{4\kappa^{2}Q^{2}}{\rho^{2}}\Bigl(l^{2}+\frac{1}{3}\Bigr)+\kappa^{2}Q^{2}\Bigg]\,u=0.\end{aligned}$
We solve this equation numerically. The numerical visualization of the wavefunctions are shown in Fig. 6.
Figure 6: The probability density $\sqrt{g}|R(r)|^{2}r^{2}$ in the radial direction. Comparison between the flat space and the Reissner–Nordström black hole background for $a=3$ and $Q=1$. Inside the inner horizon (left) and outside the outer horizon (right). The figure corresponds to the $3p$ orbital ($n=3$, $l=1$) of the hydrogen atom.
The region outside the outer horizon behaves similarly to the case of the Schwarzschild black hole. The wavefunction is strongly attracted to the outer horizon. On the other hand, we find that within the region inside the inner horizon, the wavefunction is attached to the inner horizon from the interior. This behavior is most naturally interpreted as the effect of the gravitational repulsion force observed near the core of the Reissner–Nordström black hole [27, 26]. This repulsive force is intuitively understood from the fact that $g_{00}\sim-\frac{a}{r}+\frac{Q^{2}}{r^{2}}$ in (4.1) can be identified with the gravitational potential in the weak gravity approximation. In the region $r\gg a$, the attractive force from $\frac{a}{r}$ is dominant while in the small $r$ region, the repulsive force from $\frac{Q^{2}}{r^{2}}$ overcomes the former.
@@
Figure 7: Comparison of the radial probability density in the Schwarzschild black hole and the Reissner–Nordström black hole backgrounds. The parameters are fixed at $a=4$, $Q=1.99$. The figures correspond to the $3p$ orbital ($n=3,l=1$) (left) and $4p$ orbital ($n=4,l=1$) (right) of the hydrogen atom. There appear new peaks close to the outer horizon in the Reissner–Nordström case.
@@ 4.2 Extremal limit
In the extremal limit $a^{2}=4Q^{2}$, the two horizons coincide $r_{+}=r_{-}$. The behavior of the solution (4.15) is inherited to this case. As we approach the extremum $4Q^{2}\to a^{2}$, the distance between the inner and the outer horizons decreases and it becomes eventually zero at $4Q^{2}=a^{2}$. In this limit, the gravitational repulsion takes place just inside the degenerate horizon.
Fig. 8 shows the radial probability density. Here, the ratio of the charges between the black hole and the electron is fixed. As $Q^{2}$ increases, the position of the outer horizon moves inward. As the horizon shifts toward the center, the peaks of the wave function that were previously “eaten away” re-appear. This is because increasing $Q$ with fixed $a$ results in a smaller horizon. We can refer the previous discussion in the Schwarzschild background. At the same time, as the black hole approaches the extremum, the wavefunction approaches the horizon from both inside and outside and it forms a shell of the high probability density at the degenerated horizon. This is an electron configuration unique to the black hole atom, not typically observed in the ordinary hydrogen atom. The 3D plots of the probability density $|\psi|^{2}$ are shown in Fig. 9.
Figure 8: The radial probability density $\sqrt{g}|R(r)|^{2}r^{2}$ of the $3p$ orbital ($n=3$, $l=1$) in the Reissner–Nordström case. We fix $a=6$ and vary $Q$ toward the extremal limit $Q=3$. Each figure shows $Q=1$ (left), $Q=2.5$ (middle) and $Q=3$ (right).
Figure 9: The 3D plots of the probability density $|\psi|^{2}$ for the hydrogen atom in the $3p$ orbital. The flat (left) and the extremal black hole backgrounds (right). The parameters are fixed to $a=6$, $Q=3$.
4.3 Over-extremal case
Finally, we examine the over-extremal case $4Q^{2}>a^{2}$. In this case, there are no horizons and it appears naked singularity at $r=0$. Even though there is the naked singularity, the solution $R(r)$ is regular at $r=0$. Now the repulsive term $\frac{Q^{2}}{r^{2}}$ strongly dominates over the attractive one $\frac{a}{r}$ in the over-extremal case $Q^{2}\gg a^{2}$ in the effective potential $V\sim-\frac{a}{r}+\frac{Q^{2}}{r^{2}}$. This therefore should push the wavefunction further outward from the naked singularity. The radial probability density in the over-extremal case is shown in Fig. 10.
Figure 10: Comparison of the numerical solution of the radial probability density $\sqrt{g}|R(r)|^{2}r^{2}$ in the over-extremal Reissner–Nordström background and in the flat space. The parameters are fixed to $a=6$ and $Q=4$. The point $r=0$ represents the naked singularity. The figure on the left corresponds to the $3p$ orbital ($n=3$, $l=1$) of the hydrogen atom and on the right to the $4p$ orbital ($n=4$, $l=1$).

### v2

@@ 3.2 Attractive force by potential
In the next section, we will discuss a genuine “black hole hydrogen atom,” namely, a negatively charged particle trapped by a positively charged black hole. Before discussing this truly ideal situation, we first consider a simpler setup, i.e., a particle subject to a hypothetical attractive potential in the Schwarzschild geometry. We stress that the Schwarzschild black hole cannot support a static electric potential in the strict sense. A charged source would inevitably change the Schwarzschild geometry to the one for the Reissner–Nordström solution. However, before proceeding to the rigorous treatment of the case in the next section, we first examine this setting as a simplified toy model to find the effects of the curved background on the hydrogenlike potential. This setup is justified by introducing a small probe charge to the Schwarzschild black hole and ignoring its backreaction to the geometry. We assume that there is the potential $V=-\frac{e^{2}}{r}$ producing an electriclike attractive force. We here employ the unit $4\pi\varepsilon_{0}=1$, and $e$ is the charge of the particle. We call this electron in the following. The equation in the radial direction is given by
@@
We find that $r=0$ and $r=a$ are the regular singular points. At the horizon $r=a$, the solution behaves like $R(r)\sim(r-a)^{0}$ and $(r-a)^{\frac{1}{2}}$. Then, $R(r)$ is regular but $R^{\prime}(r)$ for the latter case diverges at $r=a$.
Since the geometry is asymptotically flat, the equation coincide with the ordinary one for the hydrogen atom in $r\to\infty$. In this regime the solution should coincides with the well-known wave function. Since the energy spectrum of the hydrogen atom is determined by the asymptotic boundary condition, we can employ the energy spectrum $E_{n}=-\frac{\mu e^{4}}{2n^{2}\hbar^{2}}$ ($n=1,2,3,\ldots$) for the flat space even in the presence of the black hole. Following the prescription presented in the previous subsection, we find that the equation for the radial direction $R(x)=\mathrm{e}^{i\sigma x}H(x)$ is given by
$H^{\prime\prime}(x)+\left\{2\sigma+\frac{\frac{3}{2}}{x}+\frac{\frac{1}{2}}{x-1}\right\}H^{\prime}(x)+\frac{1}{x(x-1)}\left\{(2\sigma-a^{2}\kappa^{2}+as^{2})x-\frac{3}{2}\sigma-l(l+1)\right\}H(x)=0,$
where we have again defined $x=\frac{r}{a}$, $\sigma=\pm a\kappa$, $\kappa^{2}=\frac{-2\mu E_{n}}{\hbar^{2}}$, and $s^{2}=\frac{2\mu e^{2}}{\hbar^{2}}$. The general solution is, therefore, given by the confluent Heun’s function $H(\gamma,\delta,\varepsilon,\alpha,q,x)$, where the parameters are given by $\gamma=\frac{3}{2}$, $\delta=\frac{1}{2}$, $\varepsilon=2\sigma$, $\alpha=2\sigma-a^{2}\kappa^{2}+as^{2}$, and $q=-\frac{3}{2}\sigma-l(l+1)$.
We find that it is convenient to visualize the wave function by solving the equation numerically. To this end, we rewrite Eq. (3.15) into the following form:
@@
Figure 3: The radial probability densities of the wave function $\sqrt{g}|R(r)|^{2}r^{2}$. Comparison between the hydrogen atom in the flat space and in the black hole background cases for $a=1$. The figures correspond to the $3p$ orbital ($n=3$, $l=1$) (left), the $5d$ orbital ($n=5$, $l=2$) (middle), and the $6f$ orbital ($n=6$, $l=3$) (right) of the hydrogen atom. We will always use the parameters $\mu=e=\hbar=1$ in the following.
Figure 4: The 3D plot of the probability density $|\psi|^{2}$ for the particle under the attractive force. The flat space (left) and the black hole background (right) cases. The figures correspond to the $3d$ orbital ($n=3$, $l=2$) of the hydrogen atom. The values increase from blue to red and the event horizon $a=5$ is shown in the black color.
Even though the wave function is strongly concentrated near the horizon, the peak appears to be displaced to outside when it is slightly farther away from the horizon. This behavior can be interpreted as follows. Although electrons are gravitationally attracted toward the black hole and extremely localized in a thin shell near the horizon, the uncertainty principle causes their kinetic energy to increase when they are too close to the horizon surface. Interestingly, the localization of electrons due to the strong gravitational attraction of the black hole has the opposite effect of repulsing particles away from it. These two forces balance to minimize the energy, resulting in a finite spread and peak positions of the wave function. Paradoxically, the stronger the gravitational attraction toward the center, the farther the electrons are pushed away. The difference between the bound state and the free particle discussed in Sec. 3.1 is that the former makes standing waves and localizes sharply, while the latter does not. Then it is natural that the former is more subject to the uncertainty relation.
However, when the black hole becomes sufficiently large, the situation changes. Figure 5 shows a comparison of the radial probability densities obtained by varying the horizon position $r=a$.
(a) Radial probability density of the $3p$ orbital.
(b) Radial probability density of the $5d$ orbital.
Figure 5: Comparison of the radial probability density $\sqrt{g}\,|R(r)|^{2}r^{2}$ of (a) the $3p$ orbital ($n=3$, $l=1$) and (b) the $5d$ orbital ($n=5$, $l=2$) for different $a$ values in the Schwarzschild background with the potential $V$. Each figure shows $a=4$ (left), $a=8$ (middle), and $a=14$ (right) for the $3p$ orbital, and $a=4$ (left), $a=14$ (middle), and $a=30$ (right) for the $5d$ orbital. The red dotted line denotes the event horizon.
When the horizon position changes, the peak positions of the wave function move accordingly. By changing the horizon position and making it larger than the peak position of the probability density, the corresponding peak appears to be absorbed into the horizon. Then gravity eventually prevails over the quantum effects and electrons will inevitably localize on the horizon.
@@ 4 Black hole atoms
We now study the genuine “black hole atoms.” We introduce the single nonextremal Reissner–Nordström black hole as the positively charged atomic nucleus. In the following, we examine the quantum wave function for a negatively charged particle trapped by the black hole. This is nothing but the hydrogen atom composed of a black hole nucleus and an electron.
@@ 4.1 Black hole hydrogen atom
The nonextremal Reissner–Nordström black hole metric in the Cartesian coordinate is given by
@@
where $r^{2}=x^{2}+y^{2}+z^{2}$, and $a=\frac{2G_{\mathrm{N}}M}{c^{2}}$ and $Q^{2}=\frac{G_{\mathrm{N}}e^{2}}{c^{4}}$ are constants. The outer and the inner horizons are at $r=r_{+}$ and $r=r_{-}$, respectively, where we have defined $r_{\pm}=\frac{a\pm\sqrt{a^{2}-4Q^{2}}}{2}$. The gauge field $A_{0}=-\frac{1}{c}\frac{e}{r}$ in the solution provides the electrostatic potential. The time slice of the solution (4.1) gives a three-dimensional spatial curved space. The spatial metric is given by
@@
We also have $g=\det g_{ij}=\frac{r^{2}}{(r-r_{+})(r-r_{-})}$. Then, after some calculations, the Schrödinger equation in the $(r,\theta,\phi)$ system is given by
@@
By substituting the decomposition $\psi(r,\theta,\phi)=R(r)Y(\theta,\phi)$ in the equation, we find that the angular function $Y(\theta,\phi)$ satisfies the same equation in the Schwarzschild case, and hence it is the spherical harmonics. On the other hand, the equation in the radial direction is given by
@@
Then, $R(r)$ is regular at the horizons, but for $R(r)\sim(r-r_{\pm})^{\frac{1}{2}}$ its derivative diverges, as in the case of the Schwarzschild black hole. Now we look for bound states where $E<0$. The asymptotic consistency again results in $E_{n}=-\frac{\mu e^{4}}{2n^{2}\hbar^{2}}$. Using the same parameters as discussed before, Eq. (4.13) is rewritten as
$\begin{aligned} \displaystyle&(\rho^{2}-2\kappa a\,\rho+4\kappa^{2}Q^{2})\,\frac{\partial^{2}u}{\partial\rho^{2}}+\Bigl\{-\rho^{2}+2\rho\,(l+1+\kappa a)-\kappa a\,(4l+3)+\frac{4\kappa^{2}Q^{2}}{\rho}\,(2l+1-\rho)\Bigr\}\frac{\partial u}{\partial\rho}\\ &+\Bigg[\rho\Bigl(n-l-1-\frac{\kappa a}{2}\Bigr)+\kappa a\Bigl(2l+\frac{3}{2}\Bigr)-\left\{\frac{\kappa a\,l}{\rho}+\frac{2\kappa^{2}Q^{2}}{\rho}\right\}(2l+1)+\frac{4\kappa^{2}Q^{2}}{\rho^{2}}\Bigl(l^{2}+\frac{1}{3}\Bigr)+\kappa^{2}Q^{2}\Bigg]\,u=0,\end{aligned}$
where $\rho=2\kappa r$ and $R(r)=\rho^{l}\mathrm{e}^{-\frac{\rho}{2}}u(\rho)$ as in the case of Eq. (3.17). We solve this equation numerically. The numerical visualization of the wave functions is shown in Fig. 6.
(a) Radial probability density of the $3p$ orbital.
(b) Radial probability density of the $5d$ orbital.
Figure 6: The probability density $\sqrt{g}|R(r)|^{2}r^{2}$ in the radial direction. Comparison between the flat space and the Reissner–Nordström black hole (RN BH) background for $a=3$ and $Q=1$. Inside the inner horizon (left) and outside the outer horizon (right). The figure corresponds to (a) the $3p$ orbital ($n=3$, $l=1$) and (b) the $5d$ orbital ($n=5$, $l=2$) of the hydrogen atom.
The region outside the outer horizon behaves similarly to the case of the Schwarzschild black hole. The wave function is strongly attracted to the outer horizon. On the other hand, we find that within the region inside the inner horizon, the wave function is attached to the inner horizon from the interior. This behavior is most naturally interpreted as the effect of the gravitational repulsion force observed near the core of the Reissner–Nordström black hole [27, 26]. This repulsive force is intuitively understood from the fact that $g_{00}\sim-\frac{a}{r}+\frac{Q^{2}}{r^{2}}$ in (4.1) can be identified with the gravitational potential in the weak gravity approximation. In the region $r\gg a$, the attractive force from $\frac{a}{r}$ is dominant, while in the small $r$ region, the repulsive force from $\frac{Q^{2}}{r^{2}}$ overcomes the former.
@@
Figure 7: Comparison of the radial probability density in the Schwarzschild black hole and the Reissner–Nordström black hole backgrounds. The left figure shows the $3p$ orbital ($n=3,l=1$) with the parameters fixed at $a=5$ and $Q=2.49$, while the right figure shows the $5d$ orbital ($n=5,l=2$) with the parameters fixed at $a=10$ and $Q=4.5$. In the Reissner–Nordström case, new peaks appear near the outer horizon, indicating the characteristic influence of the electromagnetic charge on the radial distribution. There appear new peaks close to the outer horizon in the Reissner–Nordström case.
@@ 4.2 Extremal limit
In the extremal limit $a^{2}=4Q^{2}$, the two horizons coincide, $r_{+}=r_{-}$. The behavior of the solution (4.15) is inherited to this case. As we approach the extremum $4Q^{2}\to a^{2}$, the distance between the inner and the outer horizons decreases and it becomes eventually zero at $4Q^{2}=a^{2}$. In this limit, the gravitational repulsion takes place just inside the degenerate horizon.
Figure 8 shows the radial probability density. Here, the ratio of the charges between the black hole and the electron is fixed. As $Q^{2}$ increases, the position of the outer horizon moves inward. As the horizon shifts toward the center, the peaks of the wave function that were previously “eaten away” reappear. This is because increasing $Q$ with fixed $a$ results in a smaller horizon. We can refer to the previous discussion in the Schwarzschild background. At the same time, as the black hole approaches the extremum, the wave function approaches the horizon from both inside and outside and it forms a shell of the high probability density at the degenerated horizon. This is an electron configuration unique to the black hole atom, not typically observed in the ordinary hydrogen atom. The 3D plots of the probability density $|\psi|^{2}$ are shown in Fig. 9.
(a) Radial probability density of the $3p$ orbital.
(b) Radial probability density of the $5d$ orbital.
Figure 8: The radial probability density $\sqrt{g}|R(r)|^{2}r^{2}$ of (a) the $3p$ orbital ($n=3$, $l=1$) and (b) the $5d$ orbital ($n=5$, $l=2$) in the Reissner–Nordström case. We fix $a=6$ and vary $Q$ toward the extremal limit $Q=3$. Each figure shows $Q=1$ (left), $Q=2.5$ (middle), and $Q=3$ (right).
Figure 9: The 3D plots of the probability density $|\psi|^{2}$ for the hydrogen atom in the $3p$ orbital. The flat (left) and the extremal black hole backgrounds (right). The parameters are fixed to $a=6$ and $Q=3$. The values increase from blue to red and the event horizon is shown in the black color.
4.3 Overextremal case
Finally, we examine the overextremal case $4Q^{2}>a^{2}$. In this case, there are no horizons and it appears naked singularity at $r=0$. Even though there is the naked singularity, the solution $R(r)$ is regular at $r=0$. Now the repulsive term $\frac{Q^{2}}{r^{2}}$ strongly dominates over the attractive one $\frac{a}{r}$ in the overextremal case $Q^{2}\gg a^{2}$ in the effective potential $V\sim-\frac{a}{r}+\frac{Q^{2}}{r^{2}}$. This, therefore, should push the wave function further outward from the naked singularity. The radial probability density in the overextremal case is shown in Fig. 10.
Figure 10: Comparison of the numerical solution of the radial probability density $\sqrt{g}|R(r)|^{2}r^{2}$ in the overextremal Reissner–Nordström background and in the flat space. The parameters are fixed to $a=5$ and $Q=3$. The point $r=0$ represents the naked singularity. The figure on the left corresponds to the $3p$ orbital ($n=3$, $l=1$) of the hydrogen atom and that on the right to the $5d$ orbital ($n=5$, $l=2$).